OpenStudy (elleblythe):

How to solve for the 6 trigonometric function values of theta=19pi/6??

OpenStudy (anonymous):

does it help to know that this is coterminal with \[\frac{7\pi}{6}\]? find the trig functions there you got a unit circle cheat sheet?

OpenStudy (anonymous):

look at the unit circle on the last page of the attached cheat sheet find the coordinates of the point corresponding to \(\frac{7\pi}{6}\) the first coordinate is cosine, the second is sine

OpenStudy (imstuck):

It's also equal to 210 in degrees. Heres mine in degrees, not radians.

OpenStudy (anonymous):

degrees shmegrees problem is given in numbers, work with the numbers, no need to change units

OpenStudy (elleblythe):

Are these the correct answers: cos19pi/6= -√3/2 sec19pi/6= -2√3/2 sin19pi/6= -1/2 csc19pi/6= -2 tan19pi/6= -√3/3 cot19pi/6= -√3 @satellite73 @IMstuck

OpenStudy (anonymous):

looks good to me